An ethanol - propanol solution is to be separated in a plate rectification column at atmospheric pressure. The feed, of which is 50 mole percent vapor, contains 40 mole percent ethanol. 98 % of the ethanol should be recovered in the top product, which should be 95 % pure by mole. The column is provided with a total condenser and a kettle-type reboiler. If the feed rate is 200 kmol/h and the reflux ratio is set at 1.7 times its minimum value, reply the followings:

a) Draw schematically a column and show the currents and their compositions.

b) Calculate the flow rates of top and bottom products and the composition of the bottom product,

c) Write the equations of q-line, and enriching and stripping section operating lines.

d) Find the number of the ideal plates by using McCabe-Thiele Method.

e) Calculate the condenser and reboiler loads (duties) as kW.

Latent heat of vaporization of the solutions may be assumed constant at 47 000 kJ/kmol.

ethanol -propanol at 1 atm
Answer #1

Or, [=L+QF 7 L - L za =vtw lam-wynt two Gmt = Rutlam-do Ru Ru Ru= Ww=Boil upratio R feed line intercept = -to q=fraction oAdding Oth (-5) =(L-1)x+ Dot Who (-a)y qa tra F - 1924 atra 1-q 1-9 feed line ar interceptfrom mole balance and material balance all the values are calculated and also equations of all the lines are calculated. And for option (d) relative volatility is needed so that mc cabe thiele can be applied and plotted in graph to find out number of tray